A boat traveled 210 miles downstream and back. The trip downstream took 10 hours. The trip back took 70 hours. What is the speed of the boat in still water? What is the speed of the current?
so far i had 10c+70s=210m
but i don't think that's right.

Answer by checkley77(12569)

210/10=21 mph down stream.
210/70=3 mph up stream.
(21-3)/2=18/2=9 mph is the speed of the stream.
Proof:
21-9=12 mph for the speed of the boat in still water.
3+9=12 ditto.